OpenStudy (anonymous):

Suppose y varies directly with x, and y = 10 when x = –3. What direct variation equation relates x and y? What is the value of y when x = –1?

OpenStudy (aroub):

There is two ways to do this: 1) y1/x1=y2/x2 10/-3= y/-1 10(-1)=y(-3) -10=-3y y=10/3 2) y=kx 10=k (-3) k=10/-3 y=kx y=-10/3 (-1) y=10/3

OpenStudy (amistre64):

direct variation is of the form y = x ; not y= x+b

OpenStudy (anonymous):

srry, you are right..

OpenStudy (amistre64):

'sok, it happens :)